AOB

Speaker : All ( All )


Talk time : 12:40, Duration : 00:20
Previous AOB

Status summarized below.
  • Production for JetCorr
    Status: jobs were running since last week, 20% farm saturation drop due to running RUn 14 re-production (not good). Dataset added to the DataManagement system for restore as we go.
  • Sti/StiHR/StiCA evaluation
    Status: Initial evaluation showed last week. Did not have the time to see if more appeared.
  • Re-production of the picoDST - did not start yet
  • picoDST code readiness: reviews opened
    Status: Shengli took the testing from Peifeng. This is now tested and will be moved to review stage.
  • Missing files on Distributed disk reported: PWGC informed during Exchange meeting, DataManagement system catching up. Will consider this one closed (but we need to watch for users restoring entire datasets and act)
  • Space loan on data100: also addressed (moved to Xrootd) - space will be released to data production
  • Run 16 UPC stream production
    Status: jobs now running in SL17f (code updated).
  • Experiment network upgrade: done Tuesday as planned. Be ready for a phase N (or whatever it is).
    The RACF also need to move their equipment and this may be disruptive.
    Status: STAR done as scheduled. The RACF have announced that after a year of preparation for this event, they will not have the time to proceed. Total net loss for not moving: 300k$.
  • INSPIRE/ID
    Status: discussed with the management team. Full speed ahead (two actions will follow - one is to inform authors to upload the new XML author format)
No changes from the below or did  not have time to check for updates
  • NERSC/Cori production ongoing and going well - will provide an update (and bootstrap the disk space disaster we had 2 weeks aga)
  • Tickets or new
    • 3300: memory corruption in reconstruction chain found during Cori tests
    • Issues with generating picoDST > 2 GB size (no ticket / should have one)
      Status: this should be solved
  • EPD software review - will begin after the picoDST issues are closed
  • dN/dx versus dE/dx - will further check with the PWGC and request more feedback
  • Track extension - Daniel re-contacted (we dropped the ball since July, we were supposed to help Daniel with findings)
  • Tutorials - reminded the team we need to identify our best page. Juniors polled (no reaction). This effort will come soon.
  • HFT embedding - had a special meeting with the HFT team (Xin,...) to discuss this last week. Hope to provide an update in +2 months